Problem #3: Long-Run Labor Demand and Factor Substitutability Suppose there are two inputs in the production function, labor (L) and capital (K), which can be combined to produce Y units of output according to the following production function:
7. Suppose the firm can choose whatever combination of capital (K) and labor (L) it wants to produce 600 units. Suppose the price of capital is S1,000 per machine per week. What combination of inputs (K and L) will the firm use if the weekly salary of each worker is $400? 8. What f everything is same as in the previous question but the weekly salary of each worker is $300? Now what combination of inputs (K and L) will the firm use to produce its 600 units? 9. (Bonus) What is the (wage) elasticity of labor demand for this firm as the wage falls from $400 to $300?
Answer #1

7)Y = 30 K + 10 L

The above production function represent a case of perfect substitute in production factors .

marginal product of capital = MPk = dY/ dk = 30

MPk/price of capital(r) = 30 /1000 = 0.03

marginal product of labour(w) = MPl.= dY/dl = 10

MPl/price of labour = 10/400 = 0.025

since MPk/r > MPl/w

firm will use all units of capital to produce 600 units of output

total output = 600

MPk = 30

so numbest of capital units required = 600/30 = 20

(answered one question as per policy )